OpenStudy (pokemon23):

(15x+7)- 12=4 (15x+7)- 12=4 @Mathematics

OpenStudy (pokemon23):

Hi again ;)

OpenStudy (saifoo.khan):

Hey, hello there.

OpenStudy (saifoo.khan):

15x + 7 - 12 = 4 Add like terms, 15x - 5 = 4 15x = 4 + 5 15x = 9 \[x = \frac{9}{15}\] \[x = \frac 35\]